OpenStudy (anonymous):

integral 1/xln^3x dx

OpenStudy (kirbykirby):

do you mean \[\int \frac{1}{x\ln(3x)}\,dx\]

OpenStudy (kirbykirby):

o.o?

OpenStudy (kirbykirby):

It doesn't look clear in how you wrote it.

OpenStudy (anonymous):

do you mean...\[\int\limits_{}^{}(\ln3x)/xdx\]

OpenStudy (anonymous):

which one?

OpenStudy (anonymous):

Looks like \[\int\frac{dx}{x(\ln x)^3}\]

OpenStudy (anonymous):

1/x(ln^3)x

OpenStudy (anonymous):

I'm pretty sure \(\ln^3x=(\ln x)^3\)

OpenStudy (anonymous):

If that's the case, then substitute \(u=\ln x\) so that \(du=\dfrac{dx}{x}\), then \[\int\frac{dx}{x\ln^3x}=\int\frac{du}{u^3}\]